Today the current EUR to USD exchange rate is 1 EUR = 1.19 USD. According to the Bloomberg consensus estimate, the EUR to USD exchange rate in four years is forecasted to be 1 EUR = 1.31 USD. You begin with 100 USD today and will invest in a European security that provides 10% annual returns (in EUR). Using this information, answer the following six (6) questions.

1) Is the USD forecasted to appreciate or depreciate relative to EUR? a) Appreciate b) Depreciate c)Neither appreciate nor depreciate d)Both appreciate and depreciate 2) To invest in the European security you must first convert USD to EUR. How many EUR will you have when exchange your 100 USD today for EUR? 3)Given that you will invest in a European security that earns 10% annually, how many euros will you have in 4 years? Assume that you reinvest all earnings (i.e., your investment simply compounds for 4 years). 4) You live in the US so you wish to exchange EUR for USD after 4 years. How many USD will you have in 4 years after exchanging your euros from the previous problem for USD? 5)What is your total return in USD? Use decimal notation, e.g., enter 0.2345 for 23.45% 6)What is your average annual return in USD? Use decimal notation, e.g., enter 0.2345 for 23.45%. [Hint: You can use your calculators built-in financial function ("fab five") to obtain this answer.]
Answer #1

1.1 EUR will be costing more USD after four years. Hence USD will depreciate.

USD is forecasted to depreciate relative to EUR.

ANSWER:b) Depreciate

2. Spot Exchange Rate::

1EUR=1.19 USD

1USD=(1/1.19)EUR

100 USD=100*(1/1.19) EUR=84.0336 EUR

3.Future Value of Investment after 4 years =(Present Value)*((1+i)^N)

i=Interest rate=10%=0.1

N=Number of years =4

PV=84.0336 EUR

Future Value =84.0336*(1.1^4)=123.0336 EUR

4. Spot Exchange Rate after 4 years:

1EUR=1.31USD

123.0336 EUR =1.31*123.0336 USD=161.17 USD

5. Total Return:

Investment =100USD

Future Value after 4 years=161.17USD

Total Return =161.17-100=61.17 USD

Return in percentage =(61.17/100)*100%=61.17%

6. Average annual return =(61.17/4)%=15.29%